# SmartAgent: Chain-of-User-Thought for Embodied Personalized Agent in Cyber World

### v2

@@ 1 Introduction
Embodied artificial intelligence [9] is considered as a crucial stride toward Artificial General Intelligence (AGI) [10]. Powered by the recent advances in large multi-modal models, embodied agents have been built upon to behave like real humans, capable of perceiving, reasoning, and acting with their surroundings in both real and cyber worlds. The enthusiasm for deploying such humanoid capabilities is evident in various tasks, including autonomic robotics [7, 2], game AI [44, 35], smart device assistants [30, 17], and smart city [12, 42]. Many of these scenarios require embodied agents to do more than follow instructions and execute actions like emotionless robots; otherwise, they are expected to serve as personal assistants attuned to human preferences in the meantime. For instance, for smart device assistance, agents struggle to personalize responses to ambiguous user queries such as providing music recommendations, though they fully understand the operation logic of a music player. Generally speaking, a fully functional embodied agent necessitates personalized perceptual capabilities, thereby enabling a comprehensive agent-environment-user triadic perception of the world , similar to JARVIS [^1] Stands for ’Just a Rather Very Intelligent System’, a fictional artificial intelligent character in the Marvel Comics, a fictional artificial assistant created by Iron Man which can make personalized intelligent decisions for him.
@@
However, this personalized consideration is absent among the current embodied agent works, where the optimization normally relies on golden action trajectories [31, 6] or ideal task-oriented solutions [8, 22]. Although these fixed paths can effectively accomplish task goals, they can only train embodied agents to be rigid task-oriented problem solvers, overlooking the multiple valid approaches that often exist as user-oriented indicators. Furthermore, the practical environment may exhibit unpredictable behavior, such as when new functions are involved or unseen scenes are observed. In these cases, task-oriented agents are not flexible enough to even capture dynamic changes in basic task goals [23, 34], let alone discern changes in user preference. As a result, such training paradigms actually restrict the learning of user-oriented perceptual capabilities, which could be the reason for the lower performance in many embodied personalized scenarios. For example, in the daily usage of smart devices as shown in Fig. 1, there are no golden line actions to access a music pool but rather diverse paths imply different user intents [30]. That is, these normal embodied agents are designed to only be able to use the product (music App on mobile phones), which are unable to think about users’ music preferences or extract user intentions behind usage habits and collected behaviors.
@@
In this work, we propose a novel reasoning paradigm Chain-of-User-Thought (COUT), which extends embodied agents from task-only optimization to personalization-oriented optimization. We summarize COUT process as: training agents with progressive thinking from basic embodied behaviors, gradually to explicit requirements reasoning, and finally to high-level implicit personalization understanding, as shown in Fig. 1. Note that the chain of user thoughts is a general paradigm and thus can be flexibly adapted to various cyber environments and enhance personalized reasoning. However, there exist several critical challenges to achieving COUT, as follows.
@@
To support COUT research, we collect and construct the first embodied AI benchmark with explicit personalization-related evaluations, SmartSpot. It comprises five single channels and two multi-channel scenarios to simulate complex real-world environments, featuring a total of 144 episodes and over 1,400 steps. To address the above challenges, we introduce SmartAgent, the first embodied personalized agent. It takes visual observations of GUI screenshots and textual instructions as task input and generates multi-step thoughts. Specifically, the SmartAgent undergoes a two-stage training process: embodiment stage and personalization stage, as illustrated in Fig. 3.
Figure 2: The proposed embodied personalized SmartAgent.
In embodiment stage, the agent takes GUI actions and item pool screenshots as visual inputs, along with other textual contexts. After a series of encoding processes, these multimodal tokens are fed into a Perceiver model to generate specific GUI actions, regarded as Thought #1. Based on these initial GUI thoughts, a Reasoner model then serves to infer the Thought #2, identifying the user’s potential underlying requirements in a short textual output. Finally in the personalization stage, powered by the deeper-level Thought #3, the same Perceiver directly outputs the recommendation result as “Yes” or “No”.
We quantitatively evaluate SmartAgent’s functionality in comprehensive embodied and personalized sub-tasks, including GUI Grounding, Autonomous GUI Operation, Underlying Reasoning, Personalized Recommendation, and Zero-shot Reasoning, as illustrated in Fig. 2. The results indicate (i) through three-step of COUT process with efficient LoRA tuning, SmartAgent achieves the first full-stage embodied personalized reasoning; (ii) SmartAgent delivers comparable performances to state-of-the-art task-specific models on fundamental GUI Grounding and advanced Autonomous GUI Operation tasks, capable of generating accurate action commands; (iii) SmartAgent shows proficiency in reasoning explicit underlying intentions, effectively bridging surface-level operations to more implicit-level user needs; (iv) SmartAgent excels at leveraging these preceding thoughts to uncover users’ implicit items requirements; (v) SmartAgent manifests zero-shot reasoning capabilities across new channels, a hallmark of a well-established embodied agent. We also present qualitative evaluations to illustrate SmartAgent’s effectiveness and proficiency in personal assistant scenarios. Our contributions are summarized as follows:
  • We take a pioneering step to formulate a novel reasoning paradigm COUT for embodied AI, which introduces personalized concerns that have not been addressed in previous embodied AI research. To support COUT, we construct the SmartSpot evaluation benchmark as the first case environment designed to support further research.
@@ 3.1 Definition
Table 1: Datasets statistics.
@@ 3.2 Components
As illustrated in Eq. 1, we formulate the COUT process in a common cyber world case, personal assistance of smart devices, in terms of three stages of thoughts: a) Thought #1 GUI Navigation, b) Thought #2 Underlying Reasoning, and c) Thought #3 Personalized Recommendation. Thought #1 denotes the surface-level thought for basic embodied behavior on the device GUI, e.g., This action clicks a button named [Button_name]. Thought #2 denotes the deeper-level thought for explicit user preference, e.g., It seems the user needs items with [Restriction1], [Restriction2], and [Restriction3]. Thought #3 denotes the high-level thought for implicit user preference, e.g., I recommend [item#1] from the item pool.
@@ 4.1 Dataset Summary
We collect SmartSpot data from Meituan, a well-known Internet service platform in China that offers a variety of life services, including food recommendations, hotel bookings, online flight ticket sales, etc. To create more practical personal scenes, we develop SmartSpot with two scenarios: the single-channel scenario which focuses on one type of service, and the multi-channel scenario which combines two single channels. We select five of the most daily used single channels: Food, Hotel, Flight, Movie, and Medicine. For the multi-channel scenarios, we pair Flight and Food as Travel1, and Flight and Hotel as Travel2. These combinations reflect more complicated and practical situations, such as traveling to a destination and booking hotels or restaurants. The data in every scenario are GUI action episodes which contain several steps to complete an instruction. Specifically, each episode consists of three groups of steps, as shown in different colors in Tab. 1. The blue “GUI” steps denote a series of GUI actions, like entering and completing a search bar, to access an item pool, the yellow “Item Pool” step denotes the page for the found item pool, and the “RS” steps denote the details page of each item awaiting recommendations. Each step contains a GUI screenshot, a ground-truth action (with possibly a bounding box or textual value), a list of previous actions, and an episodic instruction with a corresponding underlying requirement. Finally, SmartSpot covers seven scenarios that present a wealth of personal assistant tasks, supporting more embodied personalized research. The full data statistics are illustrated in Tab. 1.
Action Space. SmartSpot provides common human-GUI interaction operations. Following AiTW [31], each action type is assigned by a action_type_id as ground truth. We detail the full action space in Sec. 8 of Appendix.
Figure 3: Two-stage training paradigm of SmartAgent.
@@ 4.2 Dataset Collection & Analysis
We curate SmartSpot following the real-life usage of this platform. The process begins with generating pairs of user instructions along with the underlying requirements. To ensure consistency, we establish intention seeds to generate them simultaneously. Specifically, we recruit annotators experienced on this platform to select 2-3 significant attributes as seeds for searching specific item pools for each channel. For example, the “[recently played]” in Fig. 1 presents one seed result, which could be other choices that users can click on. By creating multiple permutations of intention seeds, we collect a diverse batch of instruction pairs while ensuring there were no duplicates. Then, according to these instruction pairs, the annotators performed episodic GUI operations, capturing screenshots and ground-truth actions along with their bounding boxes at each step. All episodes are completed based on the annotators’ personal usage habits on the platform. In total, we gather 144 episodes with over 1,400 steps.
@@ 5.1 Overview
Given a user instruction $i\in\mathcal{I}$ to complete, the agent will navigate through an episode comprising three groups of steps: searching for an item pool, finding the item pool, and making item recommendations. At time step $t$, the agent receives a screenshot observation $o_{t}\in\mathcal{O}$. Then, the agent should takes an action $a_{t}\in\mathcal{A}$ according to its current assets $\{o_{t},i,h_{t-1}\}$, where $h_{t-1}=(o_{1},a_{1},...,o_{t-1},a_{t-1})$ is the historical actions. The chosen action signifies either a GUI command, a signal that an item pool has been found, or a recommendation result for items within that pool. We will details the action space $\mathcal{A}$ in Sec. 4
@@
Using this representation, we formulate the learning of SmartAgent as GPT-style auto-regressive approach, in line with [20]. For instance, in Fig. 1, given a smartphone screenshot and user’s instruction “Can you give me some music recommendations?”, we craft a query prompt as: “Please generate the next action according to the <screenshot>and <instruction>”. Next, we will detail the training progress, evaluation, and detailed settings.
@@ 5.2 Agent Training
Figure 4: Comparison of methods on SmartSpot. SmartAgent performs comparable and even better with a SOTA GUI specialist and general LLMs in all scenarios. Notably, in the more complex scenarios TRAVEL1 and TRAVEL2, which involve longer episodes, SmartAgent consistently shows excellent performance across all embodied and personalized metrics.
Embodied Stage. This stage aims to complete ambiguous instructions to find the item pool. The agent takes only GUI action and item pool screenshots as visual input. As stated in the Sec. 5.1, the multimodal assets first feed into the Perceiver to predict the specific embodied actions, referred to as Thought #1 in COUT. Subsequently, the Reasoner infers step-wise action thoughts and summaries as an underlying requirement as Thought #2. The Thought #2 indicates the user’s intention explicitly reflected in this stage. For instance, as illustrated in Fig. 1, for user inquiries about music, Thought #2 may include specific constraints not present in the original instruction, such as “Justin Bieber”. As a result, the underlying requirement serves as a key intermediate output, offering a clearer representation of user intentions for the next personalized stage.
@@ 5.3 Implementation Details
We choose Qwen-VL [1] as our backbone LVLM, which encodes visual inputs with a high resolution of 448*448. The training of SmartAgent starts from the continual pre-training on SeeClick base model [3] for basic GUI grounding abilities. We train SmartAgent for around 30 epochs. During training, we apply LoRA [19] to fine-tune both the visual encoder and LLM. We utilize AdamW as the optimizer, starting with a learning rate of 3e-5 and a global batch size of 14. All training is conducted on two NVIDIA A100 GPUs. More details can be found in Appendix.
@@ 6 Evaluation
We demonstrate SmartAgent’s capabilities by comprehensive evaluations of overall abilities in Sec. 6.1 and a full spectrum of sub-tasks encompassing GUI Grounding in Sec. 6.2, Autonomous GUI Operation in Sec. 6.3, Explicit & Implicit Personalized Reasoning in Sec. 6.4, and Zero-shot Reasoning in Sec. 6.5. We also report more insights in Sec. 6.6.
Metrics. Following most of the setting in [3], we compute the cross-entropy loss for Thought #1 and Thought #3 with their ground-truth actions, and semantic similarity for Thought #2 with the underlying requirement. A bounding box may be contained in ground-truth action to verify if a click action is hit. We therefore evaluate SmartAgent using the below metrics, in terms of embodied action metrics (following Mind2web [6]) for Thought #1, and personalized preference metrics for Thought #2 and Thought #3:

### v3

@@ 1 Introduction
Embodied artificial intelligence [9] is considered as a crucial stride toward Artificial General Intelligence (AGI) [10]. Powered by the recent advances in large multi-modal models, embodied agents have been built upon to behave like real humans, capable of perceiving, reasoning, and acting with their surroundings in both real and cyber worlds. The enthusiasm for deploying such humanoid capabilities is evident in various tasks, including autonomic robotics [7, 2], game AI [43, 35], smart device assistants [30, 17], and smart city [12, 41]. Many of these scenarios require embodied agents to do more than follow instructions and execute actions like emotionless robots; otherwise, they are expected to serve as personal assistants attuned to human preferences in the meantime. For instance, for smart device assistance, agents struggle to personalize responses to ambiguous user queries such as providing music recommendations, though they fully understand the operation logic of a music player. Generally speaking, a fully functional embodied agent necessitates personalized perceptual capabilities, thereby enabling a comprehensive agent-environment-user triadic perception of the world , similar to JARVIS [^1] Stands for ’Just a Rather Very Intelligent System’, a fictional artificial intelligent character in the Marvel Comics, a fictional artificial assistant created by Iron Man which can make various personalized intelligent decisions for the user.
@@
However, this personalized consideration is absent among the current embodied agent works, where the optimization normally relies on golden action trajectories [31, 6] or ideal task-oriented solutions [8, 22]. Although these fixed paths can effectively accomplish task goals, they can only train embodied agents to be rigid task-oriented problem solvers, overlooking the multiple valid approaches that often exist as user-oriented indicators. Furthermore, the practical environment may exhibit unpredictable behavior, such as when new functions are involved or unseen scenes are observed. In these cases, task-oriented agents are not flexible enough to even capture dynamic changes in basic task goals [23, 34], let alone discern changes in user preference. As a result, such training paradigms actually restrict the learning of user-oriented perceptual capabilities, which could be the reason for the lower performance in many embodied personalized scenarios. For example, in the daily usage of smart devices as shown in Figure 1, there are no golden line actions to access a music pool but rather diverse paths imply different user intents [30]. That is, these normal embodied agents are designed to only be able to use the product (music App on mobile phones), which are not able to think about users’ music preferences or extract the user intentions behind the usage habits and collected behaviors.
@@
In this work, we propose a novel reasoning paradigm Chain-of-User-Thought (COUT), which extends embodied agents from task-only optimization to personalization-oriented optimization. We summarize COUT process as: training agents with progressive thinking from basic embodied behaviors, gradually to explicit requirements reasoning, and finally to high-level implicit personalization understanding, as shown in Figure 1. Note that the chain of user thoughts is a general paradigm and thus can be flexibly adapted to various cyber environments and enhance personalized reasoning. However, there exist several critical challenges to achieving COUT, as follows:
@@
To support COUT research, we collect and construct the first embodied AI benchmark with explicit personalization-related evaluations, named SmartSpot. It comprises five single channels and two multi-channel scenarios to simulate complex real-world environments, featuring a total of 144 episodes and over 1,400 steps. To address the above challenges, we design SmartAgent, the first embodied personalized agent. It takes visual observations of GUI screenshots and textual instructions as task input and generates multi-step thoughts. Specifically, the SmartAgent undergoes a two-stage training process: embodiment stage and personalization stage, as illustrated in Figure 3. In embodiment stage, the agent takes GUI actions and item pool screenshots as visual inputs, along with other textual contexts. After a series of encoding processes, these multimodal tokens are fed into a Perceiver model to generate specific GUI actions, regarded as Thought #1. Based on these initial GUI thoughts, a Reasoner model then serves to infer the Thought #2, identifying the user’s potential underlying requirements in a short textual output. Finally in the personalization stage, powered by the deeper-level Thought #3, the same Perceiver directly outputs the recommendation result as “Yes” or “No”.
Figure 2: The full-stage embodied personalized capabilities of our proposed SmartAgent, range from basic environment cognition to advanced user personal intention reasoning.
We quantitatively evaluate SmartAgent’s functionality in comprehensive embodied and personalized sub-tasks, including GUI Grounding, Autonomous GUI Operation, Underlying Reasoning, Personalized Recommendation, and Zero-shot Reasoning, as illustrated in Figure 2. The results indicate (i) through three-step of COUT process with efficient LoRA tuning, SmartAgent achieves the first full-stage embodied personalized reasoning; (ii) SmartAgent delivers comparable performances to state-of-the-art task-specific models on fundamental GUI Grounding and advanced Autonomous GUI Operation tasks, capable of generating accurate action commands; (iii) SmartAgent shows proficiency in reasoning explicit underlying intentions, effectively bridging surface-level operations to more implicit-level user needs; (iv) SmartAgent excels at leveraging these preceding thoughts to uncover users’ implicit items requirements; (v) SmartAgent manifests zero-shot reasoning capabilities across new channels, a hallmark of a well-established embodied agent. We also present qualitative evaluations to illustrate SmartAgent’s effectiveness and proficiency in personal assistant scenarios. The contributions of this work can be summarized as follows:
  • We take a pioneering step to formulate a novel reasoning paradigm COUT for embodied AI, which introduces personalized concerns that have not been addressed in previous embodied AI research. To support COUT, we construct the SmartSpot benchmark as the first case environment for further research.
@@ 3.1 Definition
Table 1: Datasets statistics. Instruction Mean and Underlying Mean denote the mean length of user instructions and corresponding underlying requirements.
@@ 3.2 Components
As illustrated in Equation (1), we formulate the COUT process in a common cyber world case, personal assistance of smart devices, in terms of three stages of thoughts: a) Thought #1 GUI Navigation, b) Thought #2 Underlying Reasoning, and c) Thought #3 Personalized Recommendation. Thought #1 denotes the surface-level thought for basic embodied behavior on the device GUI, e.g., This action clicks a button named [Button_name]. Thought #2 denotes the deeper-level thought for explicit user preference, e.g., It seems the user needs items with [Restriction1], [Restriction2], and [Restriction3]. Thought #3 denotes the high-level thought for implicit user preference, e.g., I recommend [item#1] from the item pool.
@@ 4.1 Dataset Summary
We collect SmartSpot data from Meituan [^2] https://www.meituan.com/en-US/about-us, a well-known Internet service platform in China that offers a variety of life services, including food recommendations, hotel bookings, online flight ticket sales, etc. To create more practical personal scenes, we develop SmartSpot with two scenarios: the single-channel scenario which focuses on one type of service, and the multi-channel scenario which combines two single channels. We select five of the most daily used single channels: Food, Hotel, Flight, Movie, and Medicine. For the multi-channel scenarios, we pair Flight and Food as Travel1, and Flight and Hotel as Travel2. These combinations reflect more complicated and practical situations, such as traveling to a destination and booking hotels or restaurants. The data in every scenario are GUI action episodes which contain several steps to complete an instruction. Specifically, each episode consists of three groups of steps, as shown in different colors in Table 1. The blue “GUI” steps denote a series of GUI actions, like entering and completing a search bar, to access an item pool, the yellow “Item Pool” step denotes the page for the found item pool, and the “RS” steps denote the details page of each item awaiting recommendations. Each step contains a GUI screenshot, a ground-truth action (with possibly a bounding box or textual value), a list of previous actions, and an episodic instruction with a corresponding underlying requirement. Finally, SmartSpot covers seven scenarios that present a wealth of personal assistant tasks, supporting more embodied personalized research. The full data statistics are illustrated in Table 1.
Action Space. SmartSpot provides common human-GUI interaction operations. Following AiTW [31], each action type is assigned by a action_type_id as ground truth. The full action space is shown as follows:
  • type("text"):3. An action that types text at (x, y).
    Figure 3: Two-stage training paradigm of SmartAgent. Thought #2 serves as a key intermediate result, connecting the Embodied and Personalized stages. This summarized user underlying requirement in the GUI operations greatly narrows the scope of the item pool for the next stage of personalized reasoning and item recommendation.
  • scroll(direction): Scroll actions for the screen. Scroll up, down, left, and right are assigned with 1, 0, 8, and 9.
  • complete:10. An action of determining if has reached the item pool.
  • BACK:5. An action for returning to the previous step.
  • HOME:6. An action for returning to the homepage.
@@ 4.2 Dataset Collection & Analysis
We curate SmartSpot following the real-life usage of this platform. The process begins with generating pairs of user instructions along with the underlying requirements. To ensure consistency, we establish intention seeds to generate them simultaneously. Specifically, we recruit annotators experienced on this platform to select 2-3 significant attributes as seeds for searching specific item pools for each channel. For example, the “[recently played]” in Figure 1 presents one seed result, which could be other choices that users can click on. By creating multiple permutations of intention seeds, we collect a diverse batch of instruction pairs while ensuring there were no duplicates. Then, according to these instruction pairs, the annotators performed episodic GUI operations, capturing screenshots and ground-truth actions along with their bounding boxes at each step. All episodes are completed based on the annotators’ personal usage habits on the platform. In total, we gather 144 episodes with over 1,400 steps.
@@ 5.1 Overview
Given a user instruction $i\in\mathcal{I}$ to complete, the agent will navigate through an episode comprising three groups of steps: searching for an item pool, finding the item pool, and making item recommendations. At time step $t$, the agent receives a screenshot observation $o_{t}\in\mathcal{O}$. Then, the agent should takes an action $a_{t}\in\mathcal{A}$ according to its current assets $\{o_{t},i,h_{t-1}\}$, where $h_{t-1}=(o_{1},a_{1},...,o_{t-1},a_{t-1})$ is the historical actions. The chosen action signifies either a GUI command, a signal that an item pool has been found, or a recommendation result for items within that pool.
@@
Using this representation, we formulate the learning of SmartAgent as GPT-style auto-regressive approach, in line with [20]. For instance, in Figure 1, given a smartphone screenshot and user’s instruction “Can you give me some music recommendations?”, we craft a query prompt as: “Please generate the next action according to the <screenshot>and <instruction>”. Next, we will detail the training progress, evaluation, and detailed settings.
@@ 5.2 Agent Training
Embodied Stage. This stage aims to complete ambiguous instructions to find the item pool. The agent takes only GUI action and item pool screenshots as visual input. As stated in Section 5.1, the multimodal assets first feed into the Perceiver to predict the specific embodied actions, referred to as Thought #1 in COUT. Subsequently, the Reasoner infers step-wise action thoughts and summaries as an underlying requirement as Thought #2. The Thought #2 indicates the user’s intention explicitly reflected in this stage. For instance, as illustrated in Figure 1, for user inquiries about music, Thought #2 may include specific constraints not present in the original instruction, such as “Justin Bieber”. As a result, the underlying requirement serves as a key intermediate output, offering a clearer representation of user intentions for the next personalized stage.
@@ 5.3 Implementation Details
We choose Qwen-VL [1] as our backbone LVLM, which encodes visual inputs with a high resolution of 448*448. The training of SmartAgent starts from the continual pre-training on SeeClick base model [3] for basic GUI grounding abilities. Following the approach of [3], We intuitively present numerical coordinates as natural languages without additional tokenization. We take 8 historical actions with screenshots during training considering the GPU memory limitation. We train SmartAgent 15 epochs for both the embodied and personalized stages. All baselines are trained for 15 rounds. Results on the ScreenSpot [3] and Mind2Web [6] benchmarks are evaluated via direct testing and 10 epochs of training, as in [3]. During training, we apply LoRA [19] to fine-tune both the visual encoder and LLM. We utilize AdamW as the optimizer, starting with a learning rate of 3e-5 and a global batch size of 14. All training is conducted on two NVIDIA A100 GPUs.
@@ 6 Evaluation
We demonstrate SmartAgent’s capabilities by comprehensive evaluations of overall abilities in Section 6.1 and a full spectrum of sub-tasks encompassing GUI Grounding in Section 6.2, Autonomous GUI Operation in Section 6.3, Explicit & Implicit Personalized Reasoning in Section 6.4, and Zero-shot Reasoning in Section 6.5. We also report more insights in Section 6.6.
Metrics. Following most of the setting in [3], we compute the cross-entropy loss for Thought #1 and Thought #3 with their ground-truth actions, and semantic similarity for Thought #2 with the underlying requirement. A bounding box may be contained in ground-truth action to verify if a click action is hit. We therefore evaluate SmartAgent using the below metrics, in terms of embodied action metrics (following Mind2web) for Thought #1, and personalized preference metrics for Thought #2 and Thought #3.
Embodied action metrics:
@@
6.1 Embodied Personalized Reasoning
We first investigate the comprehensive capabilities of SmartAgent in handling embodied tasks and personalized reasoning, primarily focusing on SmartSpot for validation. Specifically, we categorize this experiment into simpler single-domain tasks and more complicated scenarios that mimic real multi-channel interactions. We exclude the MEDICINE channel to evaluate the zero-shot ability later. We select SeeClick (a widely used GUI agent backbone using pure visual observation), along with the well-known LLM foundation model, Qwen-VL series, as our baselines. We keep the original prompt reasoning setting for all general LLM and GUI Specialist baselines to show their features and deficiencies on the COUT reasoning task. Thus, only embodied action results are reported.
Results & analysis. Table 2 shows SmartAgent’s average performance across all channels. It achieved the best results in the Ele.Acc metric and ranked second in SSR, indicating strong foundational embodied perception ability. This comprehensive understanding of GUI operations supports the prediction of 71% of users’ underlying requirements, afterward leading to a 24% accuracy in item recommendations. More importantly, Figure 4 shows that SmartAgent consistently performs the best in all multi-channel scenarios, which highlights its superior reasoning capabilities in real-world applications compared to specialized GUI agents and general LLM that lack deeper personalized cognition. Qwen2-VL had unexpectedly poor results that probably were caused by highly limited tunable LoRA parameters compared to Qwen-VL under the same train epochs. Qwen2-VL will not be discussed further and the transferability of COUT across different backbone architectures warrants future investigation. Overall, these evaluations confirm SmartAgent’s comprehensive capabilities on both embodied and personalized tasks.
Figure 4: Comparison of methods on SmartSpot. SmartAgent performs comparable and even better with GUI specialist model and general LLM in all scenarios. Notably, in the more complex scenarios TRAVEL1 and TRAVEL2, which involve longer episodes, SmartAgent consistently shows excellent performance across all embodied and personalized metrics.